6. The payback period The payback method helps firms establish and identify a maximum acceptable payback period that helps inYear 2 Year o -$5,000,000 Year 1 $2,000,000 Year 3 $1,750,000 $4,250,000 Expected cash flow Cumulative cash flow ConventionalOne theoretical disadvantage of both payback methods-compared to the net present value method is that they fail to consider tAnswer in full please
Answer #1

D 1 2 3 B Year 0 (5,000,000) (5,000,000) C Year 1 2,000,000 (3,000,000) Expected cash flow Cumulative cash flow Year 2 4,250,Year 0 Year 1 Year 2 Year 3 2 Expected cash flow 3 Cumulative cash flow -5000000 =B2 2000000 =B3+C2 4250000 =C3+D2 1750000 =D